# GARG-AML against Smurfing: A Scalable and Interpretable Graph-Based Framework for Anti-Money Laundering

### v2

@@ 6 Limitations
This work has limitations related to both data availability and methodology.
@@ 6.1 Limitations in the Data
GARG-AML assumes full visibility into all transactions, which is rarely the case in practice (Altman et al., 2023; Tariq & Hassani, 2025; Dumitrescu et al., 2022). Banks typically have access only to transactions that involve their own clients, including sender and receiver account numbers and client attributes. Transactions between two external accounts remain hidden, which makes second-order neighbourhoods incomplete. This can obscure smurfing structures or lead to incorrect classification (see Figure 14). Moreover, banks cannot link multiple third-party accounts that are owned by the same individual, thus further weakening pattern recognition.
Figure 14: Illustration of smurfing over different banks, which are represented by different colours. The (i)-icon indicates that the main (blue) bank has all client information about the corresponding account. The dotted lines indicate that the bank is unaware of these transactions/accounts. In most cases, the blue bank is aware of a big part of the transaction network.
This is mainly a limitation when only one bank would be guarding against money laundering. When two banks are involved, one of these two will have a clear picture of the transaction scheme. Although the transactions go through different banks in Figures 14(a) and (b), the blue bank still has the full transaction pattern. It becomes much harder to have a full picture when more than two banks are involved.
Given these limitations, we advocate for stronger collaboration both among banks and across countries. By enabling the secure sharing of transaction information, many of the current blind spots could be substantially reduced. In such an environment, GARG-AML would be positioned to deliver significantly greater value.
@@ 6.2 Limitations in the Methodology
GARG-AML uses unweighted graph structures, ignoring transaction volume and timing. While these are simplifications, the resulting smurfing score remains interpretable and useful in downstream AML systems. Timing issues can be addressed by restricting analysis to shorter time windows, applying sliding windows, or comparing scores across temporal blocks.
@@ 7 Conclusion
This paper introduces Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), which is a novel approach for detecting smurfing by quantifying the structural similarity of nodes’ neighbourhoods to known money laundering patterns via density-based features in the adjacency matrix. We validated GARG-AML on several synthetic datasets and compared it to state-of-the-art smurfing detection methods. The results of the experiments show that GARG-AML achieves strong detection power, with state-of-the-art performance of GARG-AML scores, and excellent scalability for large networks. Unlike approaches that are based on complex optimization techniques or machine learning models, GARG-AML relies on domain expertise and straightforward density-based calculations.
While GARG-AML is not a complete solution for smurfing detection, it offers two key advantages over existing methods: (1) it produces interpretable features—either as separate densities or a single composite score—that can be seamlessly integrated into AML systems, and (2) it is computationally efficient, with parallelisable and memory-light calculations, thus enabling practical deployment in real-world settings.
Future work will explore (i) treating individual densities as features in classification models and (ii) extending the methodology to incorporate transaction timing and volume, as well as generalising it to other anomalous transaction patterns.
Additionally, money laundering schemes typically exploit cross-bank and cross-border transaction chains, which fragment the available information and obscure detection efforts. This work advocates that enhanced collaboration, both at the national level among domestic banks and at the international level across jurisdictions, is essential. Establishing secure data-sharing frameworks and cooperative monitoring mechanisms would significantly strengthen the ability to trace complex laundering networks and improve the overall effectiveness of anti-money laundering measures.
Finally, detecting smurfs is an important research area for network security and intrusion detection (Yang et al., 2009; Shadrooh & Nørvåg, 2024). Further research could analyse whether similar featurisation of the patterns is possible for denial of service (DoS) or similar attacks.
@@
Appendix A Toy Example GARG-AML Scores
We demonstrate the calculations and results as described in Section 3, using a toy example. We start with a Watts-Strogatz network containing 20 nodes, each connecting to the next two nodes (initial degree equal to four), with a rewiring probability of $20\%$. We inject a smurfing pattern, where money originates at a new node and gets transferred to an existing node in the network using three mule accounts. The full network is visualised in Figure A.1.
We select two nodes to illustrate how the GARG-AML scores are calculated. We select node 23 because its second-order neighbourhood consitutes a pure smurfing pattern. We also select node 8, since its neighbours interact a lot with each other. So, the calculations for node 8 illustrate how the scores are impacted by these interactions.
Figure A.1: The undirected network with toy example. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
@@ A.1 Undirected Calculations
We start with the pure smurfing pattern using node 23. Its second-order neighbourhood is visualised in Figure A.2. Based on this neighbourhood, we construct the adjacency matrix as follows:
@@
The elements that should be zero, and hence are not taken into account, are indicated in red. Those elements that are certainly equal to one are indicated in green.
Figure A.2: Visualisation of the undirected second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
We calculate the scores by taking the upper-left block as block 1, the upper-right as block 2, and the lower-right as block 3. Using the notation as in Equations (5)-(8) , this results in the density calculations and sizes, where the size is the number of free elements:
@@
As expected, the score that we assign to node 23 is 1. This indicates that its second-order neighbourhood forms a pure smurfing pattern.
The second-order neighbourhood of node 8 is shown in Figure A.3. We observe an interaction between node 7 and 9, which are part of the neighbours at distance one. Additionally, multiple interactions between nodes at distance two can be observed. This already indicates that the neighbourhood does not represent a pure smurfing pattern. The adjacency matrix is given in Equation (15), with the same colour scheme as before.
Figure A.3: Visualisation of the undirected second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
@@
The first row contains the node we will give a score, here node 8. The following rows represent the nodes at distance two, i.e., 3, 4, 5, 10, 11 and 15. The final rows represent the nodes at distance one, i.e., 6, 7 and 9.
@@
As indicated above, node 8 is not involved in a smurfing scheme. This is corroborated by the fact that the final score is negative.
@@ A.2 Directed Calculations
We demonstrate the calculations of the directed GARG-AML scores on the same network as before, but with directed edges. This network is illustrated in Figure A.4. We again calculate the scores for node 23 and 8.
Figure A.4: The directed network with toy example. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
The directed second-order neighbourhood for node 8 is shown in Figure A.5, and the corresponding adjacency matrix is provided in Equation (17). Note that we now have a different ordering of the nodes, using level 0 (node 23), level 1 (nodes 20, 21, 22) and level 2 (node 19), resulting in nine blocks, as defined in Section 3.3.
@@
Figure A.5: Visualisation of the directed second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
The density of the nine blocks is calculated. We will use the matrix representation as in Equation (11) to represent the densities of the blocks. For the adjacency matrix in Equation (17), this results in the following nine densities:
@@
The score for node 23 is then calculated as
@@
This again shows that the pure smurfing pattern gets a perfect score.
We do the same exercise for node 8. Here, level 0 consists of nodes 8 and 15, level 1 consists of nodes 6, 7 and 9, and level 2 consists of nodes 3, 4, 5, 10 and 11. Contrary to the second-order neighbourhood of node 23, we see on Figure A.6 that money flows in multiple directions, with nodes sending and receiving money to and from nodes across levels. This is already an indication that node 8 is not part of a smurfing scheme. The corresponding adjacency matrix is given in Equation (21).
Figure A.6: Visualisation of the directed second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
@@
We observe that the densities are much different than for the pure smurfing pattern. This translates to the following score calculation:
@@
The score for node 8 is below 0. Hence, we have no suspicion that this node is part of a smurfing scheme.
Appendix B Full Results IBM Data
In this part of the appendix, we provide the full results on the IBM datasets. These include all different patterns defined by Altman et al. (2023). The binary labels are defined using the threshold values 0.1, 0.2, 0.3, 0.5 and 0.9. The results on the HI-Small and LI-Large datasets are provided in Figure B.1 and Figure B.2, respectively.
Figure B.1: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset across all thresholds and patterns.
Figure B.2: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset across all thresholds and patterns.

### v3

@@ 6.1 Limitations in the Data
GARG-AML assumes full visibility of all transactions, which is rarely the case in practice (Altman et al., 2023; Tariq and Hassani, 2025; Dumitrescu et al., 2022). Banks only have access to transactions that involve their own clients, and the sender and receiver account numbers. Transactions between two external accounts remain hidden, which makes second-order neighbourhoods incomplete (see Figure 10). This can obscure smurfing structures or lead to incorrect classification.
Figure 10: Smurfing over multiple banks. The (i)-icon indicates that the main (blue) bank has all client information about the corresponding account. The dotted lines indicate that the bank is unaware of these transactions/accounts.
Initiatives are needed that allow to link transaction on an inter-bank and international level. However, sharing all transaction data is often not possible because of privacy and/or commercial reasons. Federated learning (McMahan et al., 2017) could provide a solution, albeit adapted to network data where links are shared among parties.
@@ 6.2 Limitations in the Methodology
There are three main limitation to the presented methodology. The first concerns the use of only one layer of smurfs. It is likely that criminal organisations utilise multiple layers of smurfs to launder their money. While multi-layered schemes exist, they impose a deterrent risk-reward trade-off: bypassing the core organization increases the risk of funds disappearing and requires greater trust in intermediaries. Furthermore, incremental smurf fees reduce the total illicit reward. Consequently, single-layer analysis already captures a meaningful portion of standard smurfing operations.
The second limitation is that relying on a single score is restrictive; it may miss complex, real-world smurfing operations (Bekkers et al., 2025; Raza et al., 2020; Leukfeldt and Jansen, 2015). An improvement would already be to use business rules on the values of the densities and sizes of the blocks. However, this requires extensive expert knowledge on diverse smurfing topologies. This is, however, outside of the scope of this paper.
The third methodological limitation is that GARG-AML uses unweighted graph structures, ignoring transaction volume and timing. While these are simplifications, the resulting smurfing score remains interpretable and useful in downstream AML systems. Timing issues can be addressed by restricting analysis to shorter time windows, applying sliding windows, or comparing scores across time.
@@ 7 Conclusion
This paper introduces Graph-Aided Risk Grading for Anti-Money Laundering (GARG-AML), which is a novel approach for detecting smurfing by quantifying the structural similarity of nodes’ neighbourhoods to known money laundering patterns via density-based features in the adjacency matrix.
We validated GARG-AML on several synthetic datasets and compared it to state-of-the-art smurfing detection methods. The results of the experiments show that GARG-AML achieves strong detection power, with state-of-the-art performance, and excellent scalability for large networks. Unlike approaches that are based on complex optimization techniques or machine learning models, GARG-AML relies on domain expertise and straightforward density-based calculations.
While GARG-AML is not a complete solution for smurfing detection, it offers two key advantages over existing methods: (1) it produces interpretable features—either as separate densities or a single composite score—that can be seamlessly integrated into AML systems, and (2) it is computationally efficient, with parallel and memory-light calculations, thus enabling practical deployment in real-world settings.
These key properties allow financial institutions to adopt our method easily and strengthen their role as capable guardians. Blocking simple smurfing tactics forces criminals into riskier, less profitable schemes. Ultimately, if the cost and danger of laundering money outweigh the rewards, regulators can effectively put criminal syndicates out of business.
Future work will explore (i) treating individual densities as features in classification models to identify distinct smurfing and money mule topologies, (ii) extending the methodology to incorporate transaction timing and volume, as well as generalising it to other anomalous transaction patterns, and (iii) constructing a federated learning framework to alleviate the limitations to the network data across banks.
Detecting smurfs is also an important research area for network security and intrusion detection (Yang et al., 2009; Shadrooh and Nørvåg, 2024). Further research could analyse whether similar featurisation of the patterns is possible for denial of service (DoS) or similar attacks.
Acknowledgments
This work was supported by the Research Foundation – Flanders (FWO) [grant numbers 1SHEN24N and G015020N] and by the BNP Paribas Fortis Chair in Fraud Analytics. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
Author Contributions
Bruno Deprez: Conceptualization, Data Curation, Methodology, Formal Analysis, Software, Visualization, Writing—Original Draft.
Bart Baesens: Conceptualization, Supervision, Writing—Review & Editing.
Tim Verdonck: Conceptualization, Supervision, Writing—Review & Editing.
Wouter Verbeke: Conceptualization, Project administration, Supervision, Writing—Review & Editing.
@@ Acknowledgments
Appendix A Illustration GARG-AML
We demonstrate GARG-AML as described in Section 3. We start with a Watts-Strogatz network containing 20 nodes, each connecting to the next two nodes (initial degree equal to four), with rewiring probability $20\%$. We inject a smurfing pattern, where money originates at a new node (23) and gets transferred to an existing node (19) in the network using three mule accounts (20, 21, 22). The full network is visualised in Figure A.1.
We calculate the GARG-AML scores for two nodes as an illustration. We select node 23 because its second-order neighbourhood constitutes a pure smurfing pattern. We also select node 8, since how the scores are impacted by the interactions of the neighbours.
Figure A.1: The undirected network. Nodes 20 to 23 are part of the infused smurfing pattern. Node 19 is the destination node.
@@ A.1 Undirected Calculations
We start with the smurfing pattern using node 23. Its second-order neighbourhood is visualised in Figure A.2. Based on this, we construct the adjacency matrix, with the blocks, as follows:
@@
The elements that should be zero are indicated in red. Those elements that are certainly equal to one are indicated in green.
We take the upper-left block as block 1, the upper-right as block 2 (can also be the lower-left, since the matrix is symmetric), and the lower-right as block 3. Using Equations (5)-(10), this results in the following:
@@
Figure A.2: Visualisation of the second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, nodes at distance two are presented in orange. All other nodes are given in gray.
The second-order neighbourhood of node 8 is shown in Figure A.3. We observe an interaction between neighbours at distance one and two, something we do not have for a pure smurfing pattern. These are reflected in the adjacency matrix in Equation (17). The first row contains the node to be scored, node 8. The following rows represent the nodes at distance two, i.e., 3, 4, 5, 10, 11 and 15. The final rows represent the nodes at distance one, i.e., 6, 7 and 9.
Figure A.3: Visualisation of the second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
@@ A.2 Directed Calculations
The directed network is illustrated in Figure A.4. We calculate the scores for node 23 and 8.
Figure A.4: The directed network. Nodes 20 to 23 are part of the infused smurfing pattern, and with node 19 the destination node.
The directed second-order neighbourhood for node 8 is shown in Figure A.5, and the corresponding adjacency matrix is provided in Equation (19). Note that we now have a different ordering of the nodes, using level 0 (node 23), level 1 (nodes 20, 21, 22) and level 2 (node 19), resulting in nine blocks.
@@
Figure A.5: Visualisation of the second-order neighbourhood of node 23 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
The density of the nine blocks is calculated for the final score. We will use the matrix representation of Equation (13) to represent the densities. This results in:
@@
The score for node 23 is calculated as
@@
We do the same for node 8. Here, level 0 consists of nodes 8 and 15, level 1 consists of nodes 6, 7 and 9, and level 2 consists of nodes 3, 4, 5, 10 and 11. Contrary to the second-order neighbourhood of node 23, we see on Figure A.6 that money flows in multiple directions, with nodes sending and receiving money to and from nodes across levels. This indicates that node 8 is not part of a smurfing scheme. The corresponding adjacency matrix is given in Equation (23).
Figure A.6: Visualisation of the second-order neighbourhood of node 8 (presented in red). Nodes at distance one are presented in blue, and nodes at distance two are presented in orange. All other nodes are given in gray.
@@
We can, indeed, see that the score for node 8 is below 0. Hence, we have no suspicion that this node is part of a smurfing scheme.
Appendix B Results Synthetic Data
The results are presented according to the injected patterns. These are provided in Figure B.1, with their corresponding ranks presented in Figure B.2.
Figure B.1 shows the distribution of AUC-ROC and AUC-PR values across the datasets. The AUC-PR values are more spread out than the AUC-ROC values. For a model with random label predictions, the AUC-ROC is always 0.5, regardless of the label distribution, whereas the AUC-PR equals the proportion of positive labels, which leads to the high variance.
Figure B.1 indicates that, on average, the undirected GARG-AML scores consistently showing great performance. This performance is boosted when combining the GARG-AML scores with the tree-based learners. In comparison, the performance of AutoAudit is less stable. Here, we set all the AUC-ROC and AUC-PR values to 0 when the method runs out of time or out of memory to allow for a full comparison of the ranks.
Figure B.2 shows the boxplots of the ranks of the methods (rank 1 means best scoring model, and rank 8 worst). We can see more clearly that combining the tree-based learners with the scores and summary statistics of the node’s neighbours score better than the other methods, even when just applying a simple decision tree.
Figure B.1: AUC-ROC and AUC-PR values of the compared methods across all synthetic datasets, with the split-on money laundering pattern.
Figure B.2: Ranks of the AUC-ROC and AUC-PR values of the compared methods across all synthetic datasets, with the split-on money laundering pattern, where rank 1 indicates the best performance and rank 8 indicates the worst performance.
Appendix C Results IBM Data
In this part of the appendix, we provide the full results on the IBM datasets. These include all patterns defined by Altman et al. (2023). The binary labels are defined using the cut-off values 0.1, 0.2, 0.3, 0.5 and 0.9. The results on the HI-Small and LI-Large datasets are provided in Figure C.1 and Figure C.2, respectively.
Figure C.1 shows the strong performance of the undirected GARG-AML score on the HI-Small. A big performance increase is observed for the tree-based methods at lower cut-offs in term of the AUC-PR. A surprising result is that the base scores seem to outperform the other models in the non-smurfing patterns as well.
Figure C.1: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the HI-Small IBM dataset.
The supervised learning methods seem to struggle with the extreme class imbalance of the LI-Large dataset. Figure C.2 shows that the best performance is obtained by the (unsupervised) base GARG-AML scores. This holds for both the directed and undirected network. This illustrates the power of GARG-AML when faced with a realistic low ratio of labelled money laundering cases.
Figure C.2: The AUC-ROC (solid bars) and AUC-PR (hatched bars) results on the LI-Large IBM dataset.